OpenStudy (unklerhaukus):

\[\{2,1,3,4,7,11,...,...,...\}\]

OpenStudy (phoenixfire):

I don't know how 2 became 1, but the rest follow a pattern.

OpenStudy (unklerhaukus):

what is the pattern ,

OpenStudy (phoenixfire):

\[x_{i+1}=x_i+x_{i-1}\]

OpenStudy (phoenixfire):

Where \[i\ge 1\]

OpenStudy (unklerhaukus):

so what are then next few terms in the sequence

OpenStudy (phoenixfire):

11+7=18 18+11=29 29+18=47

OpenStudy (anonymous):